OpenStudy (anonymous):

Simultaneous equation: 4s-3t=4 3s-5t=14 Please can you show your working out

OpenStudy (anonymous):

Make the 's' the same. 4s - 3t = 4 (x3) 3s - 5t = 14 (x4) 12s - 9t = 12 12s - 20t = 56 subtract one from the other -9t = 12 -20t = 56 - 11t = -44 t = -11 sub back into one of the originals 4s - 3t = 4 4s - (-33) = 4 4s + 33 = 4 4s = -29 s = -7.25 hope this helps :}